8. 9. 10. C) Hint: Try a trig substitution. 1 So (x2 + 92dx
Answer #1

Janajzane Giveni | Cugat du (x2+932 lat n = 3tant if v-o, tao du= 3 Sec?t-dt if u=3 t = Tily 3secz do Scelanesia I gue tantew du Lt Sin 21-) Tely 2 10% 28 () sin (15/2) = 1 108 (1) น (1) --- 11 L+2 216 3 du โt2 (หน) 216